substitution 3m + 4k = 14
2m + k = 1

with steps too

Answer by unlockmath(1688)

Hello,
With these two 3m + 4k = 14
2m + k = 1
Let's work on the second equation by getting k by itself:
k=1-2m Now we can substitute this into the first:
3m + 4(1-2m)=14
rewrite this as:
3m+4-8m=14
Now combine like terms, subtract 4 and divide by -5 to get:
m=-2
Plug this into either equation and we get:
k=5
Make sense?
RJ
www.math-unlock.com